OpenStudy (anonymous):

If the volume of a gas is 24 cubic centimeters when the pressure is 16 pounds, what is the volume under a pressure of 30 pounds?

OpenStudy (anonymous):

sqrt(8)/sqrt(12) = x/75 solve for x = 61 lbs notice that when it said inversely, the new force is smaller than old force because new distance is bigger than old distance. 2) 2/6 = x/3000 x = $1000 notice that when it said inversely, the new $ is smaller than old $ because new year is bigger than old year. 3) directly, the more gas you use, the more it costs. 4) what's the question? 5) directly, more men can build more houses 6) 16/30 = x/24 x = 12.8 cubic centimeter notice that when it said inversely, the new volume is smaller than old volume because new pressure is bigger than old pressure. 7) 132/x = 32/4 x = 16.5 notice that when it said inversely, the new x is smaller than old x because new y is bigger than old y.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can i have a medal ? :)

OpenStudy (wolf1728):

It looks as if you would need to use Boyle's Law

OpenStudy (anonymous):

duuuude u must be genius!! 348 fans!!!

OpenStudy (wolf1728):

P1 * V1 = P2 * V2 V2 = P1 * V1 / P2 V2 = 16 * 24 / 30 V2 = 12.8